OpenStudy (anonymous):

Solve 2(x + 1) + 4 = 8.

OpenStudy (anonymous):

x = 1

OpenStudy (tester97):

\[2(x+1)+4=8\]Distribute the 2\[2x+2+4=8\]Add like terms \[2x+6=8\]Subtract 6 from both sides\[2x=2\]Divide by 2 on both sidess \[x=1\] \(\Huge\cal\color{pink}{Viola!}\)